OpenStudy (anonymous):

Solve 5(2x + 1) = 12x – 1. You must show all work.

OpenStudy (anonymous):

solve it distributively

OpenStudy (anonymous):

10x + 5 = 12x - 1

OpenStudy (diyadiya):

first distribute 5(2x+1) = 12x -1 5*2x +5*1 =12x -1 10x + 5 =12x -1 can you do now?

OpenStudy (anonymous):

then you solve for x